OpenStudy (anonymous):

use the given data to find the minimum sample size required to estimate a population proportion or percentage. Margin of error: seven percentage points, confidence level 95%: from prior study. phat is estimated by the decimal equivalent of 48%. n=?

OpenStudy (amistre64):

remember how to find E? this is just solving that for n

OpenStudy (amistre64):

\[E=z\sqrt{\frac{pq}{n}}\] \[\sqrt n=\frac zE \sqrt{pq}\] \[ n=\left(\frac zE \right)^2~pq\]

OpenStudy (amistre64):

\[n=\left(\frac {1.96}{.07} \right)^2~.48(.52)\]

OpenStudy (amistre64):

and round up ...

OpenStudy (anonymous):

would it be 2018

OpenStudy (amistre64):

no

OpenStudy (anonymous):

so 196

OpenStudy (amistre64):

id go with 196 ... assuming of course i read this correctly :)

OpenStudy (anonymous):

since i must round to the nearest integer?

OpenStudy (anonymous):

thank you for the help

OpenStudy (amistre64):

youre welcome